A father is thrice as old as his son . Five years Ago , he was four times as old. How old is the father?
Answer by amoresroy(361)

A father is thrice as old as his son . Five years Ago , he was four times as old. How old is the father?
Let x = age of the son
3x = age of the father
3x-5 = 4(x-5)
3x-5 = 4x -20
20-5 = 4x -3x
x = 15
3x = 3(15) = 45
The father is 45 years old.